OpenStudy (anonymous):

If the only given information of a question is the density, temperature and pressure of a mixture of oxygen and krypton, how would you find the mole percentage of krypton?

OpenStudy (anonymous):

Well, I only got up to finding the molar mass of the mixture: using MM=dRT/P

OpenStudy (anonymous):

I do not remember the exact numbers of each given information, I only recalls getting 51.33g for the molar mass of the mixture.

OpenStudy (wolfe8):

So the density, temp, and pressure are all for the mixture?

OpenStudy (anonymous):

yes

OpenStudy (wolfe8):

Not separate?

OpenStudy (lastdaywork):

@yociyoci Did the question provide density or vapour density ??

OpenStudy (anonymous):

density

OpenStudy (lastdaywork):

You already found the molar mass of the mixture. Assume the ration of moles of oxygen and krypton = x:1

OpenStudy (lastdaywork):

This will give you an equation - (molar mass of oxygen)x + (molar mass of krypton) = molar mass of mixture I think, the rest would be obvious :)

OpenStudy (lastdaywork):

@yociyoci Does that ^^ helps ??

OpenStudy (anonymous):

Umm.... I don't really understand...

OpenStudy (lastdaywork):

Oops..my equation was terribly wrong :'(

OpenStudy (lastdaywork):

Consider this - Molar mass of a sample = (total mass) / (total number of moles) = (16x + 83.8) / (x + 1) It ^^ is analogous to the "center of mass" concept of physics XD

OpenStudy (anonymous):

(32.0x+83.8)/(x+1) ?

OpenStudy (lastdaywork):

Yea..32 (I forgot it is O2)

OpenStudy (anonymous):

MM=(32.0x+83.8)/(x+1) 51.33g=(32.0x+83.8)/(x+1) 51.33x+51.33=32.0x+83.3 19.33x=32.47 x=1.68 mol O2

OpenStudy (anonymous):

What would be the next step?

OpenStudy (lastdaywork):

'x' is just a ratio of moles of O2 : moles of Kr You can use it to find the mole fraction or mole percentage..try yourself and I'll check it :)

OpenStudy (anonymous):

1.68 mol O2 : 1 mol Kr 1/(1.68+1) x 100%= 37.3% Kr

OpenStudy (lastdaywork):

:)

OpenStudy (anonymous):

Thank you!!! Thanks for helping me figure this out~ =D

OpenStudy (lastdaywork):

You should also thank @wolfe8 for staying on OS 24/7 I always check his last few posts to find who needs help.. :D

OpenStudy (wolfe8):

Huh? I barely helped today. I went out.

OpenStudy (lastdaywork):

You at-least filter out those who need help from those who just want an answer (and of-course those who simply don't bother to close their question)

OpenStudy (anonymous):

Thanks to both of you guys!! I really appreciate you guys staying with me for this hour to work this out! Especially @wolfe8 for helping me with chemistry these couple for days! Last night, we spent hours trying to figure out a solution, although it's very late, he stayed with me until I figured out the problem!!!

OpenStudy (anonymous):

I really appreciate this!! Thank you so much~ =D

OpenStudy (lastdaywork):

:)

OpenStudy (wolfe8):

Eh you're welcome >w< I've been waiting here until you solve it while my brain is asleep

OpenStudy (lastdaywork):

Take some rest Mr Zombie :P

OpenStudy (anonymous):

Without you guys, I wouldn't be able to do so well on my test!

OpenStudy (anonymous):

I had my test today and it wasn't that bad, I managed it pretty well!! =D

OpenStudy (lastdaywork):

That's nice to know...I have a test tomorrow :'(

OpenStudy (anonymous):

Good luck!!! You'll do fine~

OpenStudy (lastdaywork):

Thanks :) Bye..

OpenStudy (anonymous):

Goodnight.

OpenStudy (wolfe8):

Good luck and sleep well.